# No Regrets: Investigating and Improving Regret Approximations for Curriculum Discovery

### v1

@@ Abstract
What data or environments to use for training to improve downstream performance is a longstanding and very topical question in reinforcement learning. In particular, Unsupervised Environment Design (UED) methods have gained recent attention as their adaptive curricula enable agents to be robust to in- and out-of-distribution tasks. We ask to what extent these methods are themselves robust when applied to a novel setting, closely inspired by a real-world robotics problem. Surprisingly, we find that the state-of-the-art UED methods either do not improve upon the naïve baseline of Domain Randomisation (DR), or require substantial hyperparameter tuning to do so. Our analysis shows that this is due to their underlying scoring functions failing to predict intuitive measures of “learnability”, i.e., in finding the settings that the agent sometimes solves, but not always. Based on this, we instead directly train on levels with high learnability and find that this simple and intuitive approach outperforms UED methods and DR in several binary-outcome environments, including on our domain and the standard UED domain of Minigrid. We further introduce a new adversarial evaluation procedure for directly measuring robustness, closely mirroring the conditional value at risk (CVaR). We open-source all our code and present visualisations of final policies here: https://github.com/amacrutherford/sampling-for-learnability.
@@ 1 Introduction
Curriculum discovery—automatically generating environments for reinforcement learning (RL) agents to train on—remains a longstanding and active area of research [1, 2]. Automated curriculum learning (ACL) methods offer the potential to generate diverse environments, leading to the development of more general and robust agents. Recently, a class of methods under the umbrella of Unsupervised Environment Design (UED) has gained popularity, owing to their theoretical guarantees of robustness and empirical improvements in out-of-distribution generalisation [3, 4, 5, 6]. However, many of these methods are evaluated using a narrow set of environments, with a strong focus on gridworld mazes [7, 8]. This issue is compounded by the advent of GPU-accelerated environments and algorithm implementations [9, 10, 11, 12], leading to a lack of support for a wide range of challenging benchmark environments in current open-source UED codebases [7, 8]. In this paper, we explore how well these UED methods generalise to a different task, specifically JaxNav, a continuous single- and multi-robot navigation task in a partially observable setting modelled via range-based observations. JaxNav modifies the gridworld setting, making it closer to a real-world robotics application.
Perhaps surprisingly, we find that current state-of-the-art (SoTA) UED approaches either fail to outperform the naïve baseline of Domain Randomisation (DR) or require significant hyperparameter tuning to do so. Our analysis reveals that the scoring functions used by these methods to prioritise environment configurations for training fail to predict intuitive notions of learnability. In binary outcome settings, these “learnable” configurations are those where an agent’s success rate is neither 100% nor 0%, providing a clear learning signal. Further evidence that the scoring functions are the key weakness of SoTA UED approaches is that true regret vastly outperforms DR on the domains where we can compute it.
Based on these results we develop a different scoring mechanism which estimates learnability by rolling out the current policy on uniformly sampled maps and creating a buffer based on maps that are solved sometimes, but not always. We find that this simple and intuitive approach outperforms Domain Randomisation [13], Prioritised Level Replay [4, 5] and ACCEL [6] on four challenging environments, including our single- and multi-agent navigation domain, Xland-Minigrid and the commonly used Minigrid environment.
To clearly demonstrate the supremacy of this approach, we also introduce a new evaluation protocol for curriculum methods, designed to explicitly test robustness. Rather than evaluating on a set of arbitrary hand-designed set of environment configurations, as is done for current SoTA UED methods, our protocol computes a risk metric on the performance of the method, by evaluating its performance in the worst $\alpha\%$ of a newly sampled set of environments.
@@
    We illustrate the inefficacy of current UED methods on a robot navigation task environment.
@@ 2.1 Reinforcement Learning & UPOMDPs
We model the reinforcement learning problem as an underspecified partially observable Markov decision process (UPOMDP) [3], denoted by $\mathcal{M}=\langle A,O,\Theta,S,\mathcal{T},\mathcal{I},\mathcal{R},\gamma\rangle$. Here, $A$, $S$, and $O$ represent the action, state, and observation spaces, respectively. The agent receives an observation $o$ (without directly knowing the true state $s$) and selects an action, which results in a transition to a new state, a new observation, and an associated reward.
$\Theta$ represents the set of possible parameters, where each $\theta\in\Theta$ defines a specific level. Each $\theta$ corresponds to a particular instantiation of the POMDP, with an associated transition function $P_{\theta}:S\times A\to\Delta(S)$ and an observation function $\mathcal{I}_{\theta}:S\to O$.
@@ 2.2 Unsupervised Environment Design (UED)
Regret-based UED methods offer an alternative by generating levels that maximise the agent’s regret. Here, the regret of a policy $\pi$ on a level $\theta$ is defined as the difference between the policy’s discounted return on $\theta$ and the optimal return achievable on that level, expressed as $U(\pi\theta^{\star})-U(\pi)$, where $U$ denotes the policy’s discounted return and $\pi_{\theta}^{\star}$ is the optimal policy on $\theta$. This approach deprioritises unsolvable levels and should, in theory, generate levels on which the agent can continue to improve. Moreover, using regret as the adversary’s objective provides additional theoretical benefits. At the convergence of the two-player zero-sum game, the agent’s policy enjoys minimax regret robustness, meaning that the maximum regret over the entire level space $\Theta$ is bounded [3].
@@
Positive Value Loss (PVL). This score function approximates regret as the average of the value loss across all timesteps where it is positive. When using GAE [15] (as is done in PPO [16]), the PVL can be written as follows:
@@ 4.1 Defining Learnability
Similarly to the “Goals of Intermediate Difficulty” objective proposed by Florensa et al. [1], we desire agents to learn on levels that they can solve sometimes but have not yet mastered. Such levels hold the greatest source of possible improvement for an agent’s policy and so a successful autocurricula method must be able to find these. Indeed, given a success rate (i.e., the fraction of times the agent solves the level) of $p$ on a given level, we define learnability to be $p\cdot(1-p)$.
Intuitively, we justify our definition as follows (in a goal-based setting where there is only a nonzero reward for reaching the goal):
@@
    $p$ represents how likely the agent is to obtain positive learning experiences from a level.
@@
    $1-p$ represents the maximum potential improvement the agent can make on that level.
@@
    Multiplying these yields (probability of improvement) * (improvement potential), i.e., expected improvement.
@@ 5 Experimental Setup
Here, we outline the domains and methods used for evaluation along with our adversarial evaluation protocol. Rather than taking the common approach of reporting average performance on a set of hand designed levels (which by their very nature are arbitrary), we sample a large set of levels and examine each method’s performance on their worst-case levels. This directly targets the tails of the level distribution and as such is a superior measure of robustness. We additionally report the comparative performance of methods on the sampled set to determine the degree to which one method dominates another in terms of solvability.
@@ 7.1 Single-Agent JaxNav
Figure 3(a) shows the CVaR results on single-agent JaxNav. We find that optimising for learnability—as our method does—results in superior robustness over a wide range of $\alpha$ values, despite all methods performing similarly with $\alpha=100\%$ (which amounts to expected success rate over the entire distribution). In this plot, we also plot the results of an oracle method named Perfect Regret. This uses the same procedure as SFL but with the score function: $1-p(\text{success})$. Importantly (and different to all other methods), this method only samples solvable levels, so this metric corresponds closely to regret. While not shown here, using the same metric with unrestricted level sampling—which is a more realistic setting—performs poorly due to it prioritising unsolvable levels. In Figure 4, we perform pairwise comparisons of each baseline against our approach. We find that there are a large number of environments that all methods solve (as illustrated bright top-right corner). However, the bottom-right of the heatmaps are generally brighter than the top-left, indicating that SFL performs better in general. Overall, SFL’s superiority, and Perfect Regret’s strong performance, indicates that the flawed approximations of regret are responsible for UED’s lack of performance. We provide further evidence for this claim by running ablations in Section I.2, where we use learnability as a score function within PLR and ACCEL.
@@ 8 Related Work
Unsupervised Environment Design (UED) has emerged as a prominent method in curriculum learning for reinforcement learning (RL), promising robust agent training through adaptive curricula. Early works focused on learning potential, where the improvement in an agent’s performance determined the choice of training levels [29, 1, 30, 2]. However, robustness-oriented methods such as adversarial minimax introduced the notion of training on levels that minimise agent performance, though these often resulted in infeasible scenarios offering no learning benefits [14, 31, 3]. Minimax regret, a more refined robustness approach, alleviates some issues of adversarial sampling by ensuring the chosen levels are learnable [3, 5, 6]. However, recent work [32] demonstrated that even when regret can be exactly computed, it does not always correspond to learnability, and this mismatch can lead to stagnation during training. Our work extends this line of research by introducing a new scoring mechanism that estimates expected improvement, targeting environments with a positive but not perfect solve rate. Unlike existing minimax regret methods, our approach directly optimises for learnability, instead of using an imperfect proxy for regret, leading to more effective training on our domains. Robust RL methods have the goal of improving an agent’s robustness to environmental disturbances, and worst-case environment dynamics [33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45]. However, these methods generally consider continuous perturbations instead of a mix of discrete and continuous environment settings. Furthermore, these methods tend to be overly conservative and prioritise unsolvable levels.
@@ 9 Discussion and Limitations
In this work we only consider deterministic, binary outcome domains and due to the nature of the learnability score, SFL is only applicable to such settings. In a non-binary-outcome domain, we could potentially reuse the intuition that $p(1-p)$ is the variance of a Bernoulli distribution; in a continuous domain, an analogous metric would be the variance of rewards obtained by playing the same level multiple times. Furthermore, our implementation of SFL is in JAX but the method is general. However, one must take the cost of SFL’s additional environment rollouts into account when considering implementing our algorithm; we chose JAX because its speed and parallelisation significantly alleviates this constraint. Next, while most current SoTA UED methods, including SFL, randomly generate and curate levels, this approach may become infeasible when the environment space is vast, as random generation may have a very low likelihood of generating valid levels. Finally, while JaxNav does have deterministic dynamics, Fan et al. [46] had success transferring an RL-based multi-robot navigation policy from a simulator of identical fidelity to the real world, suggesting this should be equally possible with JaxNav.
@@ 10 Conclusion
In this paper we investigate to what extent the currently popular branch of curriculum discovery methods based on scoring functions that approximate minimax regret can be applied to multi-robot navigation. We find that the poor quality of the scoring functions used results in poor performance of UED methods, and propose a different metric based on an intuitive notion of learnability, which improves the robustness of the final policies. We also introduce a new evaluation protocol, reporting a risk measure on performance over the $\alpha\%$ worst-case (but in principle solvable) levels for each method. We hope that our findings provide inspiration for future work on more general and domain agnostic scoring functions and we open-source all of our code to facilitate this process. Ultimately, we believe that this work is a stepping stone towards bridging the gap between popular testbeds for UED and real-world applications, such as multi-robot navigation, and serves an important reality check for this budding subfield of machine learning.
@@ Appendix F Extended analysis of UED Score Functions
An extension of Section 4.2, Figure 12 illustrates the correlation between mean reward and the two most popular regret scores, MaxMC and PVL. These graphs illustrate that the trends seen for success rate also hold for episodic reward, this is expected for our environment as the two are strongly correlated. In Figure 13 we conduct the same analysis for the L1 Loss Score.

### v3

@@ Abstract
What data or environments to use for training to improve downstream performance is a longstanding and very topical question in reinforcement learning. In particular, Unsupervised Environment Design (UED) methods have gained recent attention as their adaptive curricula promise to enable agents to be robust to in- and out-of-distribution tasks. This work investigates how existing UED methods select training environments, focusing on task prioritisation metrics. Surprisingly, despite methods aiming to maximise regret in theory, the practical approximations do not correlate with regret but with success rate. As a result, a significant portion of an agent’s experience comes from environments it has already mastered, offering little to no contribution toward enhancing its abilities. Put differently, current methods fail to predict intuitive measures of “learnability.” Specifically, they are unable to consistently identify those scenarios that the agent can sometimes solve, but not always. Based on our analysis, we develop a method that directly trains on scenarios with high learnability. This simple and intuitive approach outperforms existing UED methods in several binary-outcome environments, including the standard domain of Minigrid and a novel setting closely inspired by a real-world robotics problem. We further introduce a new adversarial evaluation procedure for directly measuring robustness, closely mirroring the conditional value at risk (CVaR). We open-source all our code and present visualisations of final policies here: https://github.com/amacrutherford/sampling-for-learnability.
@@ 1 Introduction
Curriculum discovery—automatically generating environments for reinforcement learning (RL) agents to train on—remains a longstanding and active area of research [1, 2]. Automated curriculum learning (ACL) methods offer the potential to generate diverse environments, leading to the development of more general and robust agents. Recently, a class of methods under the umbrella of Unsupervised Environment Design (UED) has gained popularity, owing to their theoretical guarantees of robustness and empirical improvements in out-of-distribution generalisation [3, 4, 5].
Currently, the most popular and empirically successful subfield of UED develops methods that aim to maximise regret—the difference in performance between an optimal agent and the current agent [3, 4, 5, 6, 7, 8]. However, computing regret is intractable in all but the simplest tasks, forcing practical methods to instead approximate it [9, 4]. While the prevailing assumption has been that these approximations are faithful, we investigate this further and find that this is not the case: specifically, these scoring functions correlate with success rate rather than regret. This means that these methods tend to prioritise tasks that the agent can already solve, leading to much of the collected experience not contributing to learning an improved policy.
While we find that regret performs well in settings where we can compute it—confirming that the underlying theory is sound—we show that the common approximations are unreliable. Therefore, we focus on a different scoring mechanism which instead prioritises levels that provide a clear learning signal to the agent. More specifically, these levels are those where the agent’s success rate is neither 100% nor 0%, i.e., levels that are neither too difficult nor too easy [1, 10].
Using this scoring function, we develop Sampling For Learnability (SFL), a method that estimates learnability by rolling out the current policy on randomly sampled levels and selecting those that the agent solves sometimes, but not always. We find that this simple and intuitive approach outperforms DR [11], Prioritised Level Replay [9, 4] and ACCEL [5] on four challenging environments, including our novel single- and multi-agent robotic navigation domain, Xland-Minigrid [12] and Minigrid [13].
To truly put our method to the test, we develop a new, more rigorous robustness evaluation protocol for ACL, and demonstrate that SFL significantly outperforms all other methods. Rather than evaluating on a set of arbitrary hand-designed environment configurations, our protocol computes a risk metric on the performance of the method, by evaluating its performance in the worst $\alpha\%$ of a newly sampled set of environments.
@@
    We illustrate the inefficacy of current UED methods on several domains, including a novel robot navigation environment.
@@ 2.1 Reinforcement Learning & UPOMDPs
We model the reinforcement learning problem as an underspecified partially observable Markov decision process (UPOMDP) [3], denoted by $\mathcal{M}=\langle A,O,\Theta,S,\mathcal{T},\mathcal{I},\mathcal{R},\gamma\rangle$. Here, $A$, $S$, and $O$ represent the action, state, and observation spaces, respectively. The agent receives an observation $o$ (without directly knowing the true state $s$) and selects an action, which results in a transition to a new state, a new observation, and an associated reward. $\Theta$ represents the set of possible parameters, where each $\theta\in\Theta$ defines a specific level. Each $\theta$ corresponds to a particular instantiation of the POMDP, with an associated transition function $P_{\theta}:S\times A\to\Delta(S)$ and an observation function $\mathcal{I}_{\theta}:S\to O$.
@@ 2.2 Unsupervised Environment Design (UED)
Regret-based UED methods offer an alternative by generating levels that maximise the agent’s regret. Here, the regret of a policy $\pi$ on a level $\theta$ is defined as the difference between the policy’s discounted return on $\theta$ and the optimal return achievable on that level, expressed as $U(\pi_{\theta}^{\star})-U(\pi)$, where $U$ denotes the policy’s discounted return and $\pi_{\theta}^{\star}$ is the optimal policy on $\theta$. This approach deprioritises unsolvable levels and should, in theory, generate levels on which the agent can continue to improve. Moreover, using regret as the adversary’s objective provides additional theoretical benefits. At the convergence of the two-player zero-sum game, the agent’s policy enjoys minimax regret robustness, meaning that the maximum regret over the entire level space $\Theta$ is bounded [3].
@@
Positive Value Loss (PVL). PVL approximates regret as the average of the value loss across all timesteps where it is positive. When using GAE [15] (as in PPO [16]), PVL can be written as follows:
@@ 4.1 Defining Learnability
Similarly to the Goals of Intermediate Difficulty objective proposed by Florensa et al. [1] and the ProCuRL curriculum strategy proposed by Tzannetos et al. [10], we desire agents to learn on levels that they can solve sometimes but have not yet mastered. Such levels hold the greatest source of possible improvement for an agent’s policy and so a successful autocurricula method must be able to find these. Indeed, given a success rate (i.e., the fraction of times the agent solves the level) of $p$ on a given level, we follow Tzannetos et al. [10] and define learnability to be $p\cdot(1-p)$. In a goal-based setting where there is only a nonzero reward for reaching the goal, we justify this definition as follows:
@@
    $p$ represents how likely the agent is to obtain positive learning experiences from a level, while $1-p$ represents the maximum potential improvement the agent can make on that level. Multiplying these yields (probability of improvement) $\cdot$ (improvement potential), i.e., expected improvement.
@@
    Tzannetos et al. [10] derive this definition for two specific, simple, learning settings and show that at each training step, selecting for the highest learnability is equivalent to greedily optimising the agent’s expected improvement in its training objective.
@@ 5 Experimental Setup
We now outline the domains used along with our adversarial evaluation protocol. Rather than taking the common approach of reporting average performance on a set of hand designed levels (which by their very nature are arbitrary), we sample a large set of levels and examine each method’s performance on their worst-case levels. This directly targets the tails of the level distribution and as such is a superior measure of robustness. We further report the comparative performance of methods on the sampled set to determine the degree to which one method dominates another.
@@ 7.1 Single-Agent JaxNav
Figure 3(a) shows the CVaR results on single-agent JaxNav. We find that optimising for learnability—as our method does—results in superior robustness over a wide range of $\alpha$ values, despite all methods performing similarly with $\alpha=100\%$ (which amounts to expected success rate over the entire distribution). In this plot, we also plot the results of an oracle method named Perfect Regret. This uses the same procedure as SFL but with the score function: $1-p(\text{success})$. Importantly (and different to all other methods), this method only samples solvable levels, so this metric corresponds closely to regret. While not shown here, using the same metric with unrestricted level sampling—which is a more realistic setting—performs poorly due to it prioritising unsolvable levels. In Figure 4, we perform pairwise comparisons of each baseline against our approach. We find that there are a large number of environments that all methods solve (the bright top-right corner). However, the bottom-right is generally brighter than the top-left, indicating that SFL performs better in general. Overall, SFL’s superiority, and Perfect Regret’s strong performance, indicates that the flawed approximations of regret are responsible for UED’s lack of performance. We provide further evidence for this claim in Section I.2, where we use learnability as a score function within PLR and ACCEL.
@@ 8 Related Work
Unsupervised Environment Design (UED) has emerged as a prominent method in the ACL field, promising robust agent training through adaptive curricula. Early works focused on learning potential, where the improvement in an agent’s performance determined the choice of training levels [31, 1, 32, 2]. However, robustness-oriented methods such as adversarial minimax introduced the notion of training on levels that minimise agent performance, though these often resulted in infeasible scenarios offering no learning benefits [14, 33, 3]. Minimax regret (MMR), a more refined robustness approach, alleviates some of these issues by ensuring the chosen levels are learnable [3, 4, 5]. However, recent work [8] demonstrated that even true regret does not always correspond to learnability, and this mismatch can lead to stagnation during training. Our work extends this line of research by utilising a scoring mechanism that estimates expected improvement, targeting environments with a positive but not perfect solve rate. Unlike existing MMR methods, our approach directly optimises for learnability, instead of using an imperfect proxy for regret, leading to more effective training on our domains.
Relatedly, Tzannetos et al. [10] introduce ProCuRL, which uses a similar learnability score as SFL (and further introduce an approximation to the solve rate $p$ using the agent’s value function). However, their problem setting is distinct from ours as they assume only a limited fixed pool of tasks are utilised during training, with the goal of improving an agent’s performance over a uniform distribution over this pool. We, instead, consider the standard UED setting where we can sample an effectively unbounded number of tasks from some large distribution $\Theta$, with the goal of achieving an agent that is robust to worst-case settings and can generalise to unseen problems. Following on from this, Tzannetos et al. [34] extend ProCuRL to the setting where the target distribution of tasks is given, and they take into account both how learnable the selected task is, as well as how correlated it is with learnable tasks from the target distribution. While this approach outperforms the original method, it tackles a different problem to SFL and UED in general, i.e., where the target distribution is known.
Finally, robust RL methods have the goal of improving an agent’s robustness to environmental disturbances, and worst-case environment dynamics [35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47]. However, these methods generally consider continuous perturbations instead of a mix of discrete and continuous environment settings. Furthermore, these methods tend to be overly conservative and prioritise unsolvable levels.
@@ 9 Discussion and Limitations
In this work we only consider deterministic, binary outcome domains and due to the nature of the learnability score, SFL is only applicable to such settings. In other domains, we could potentially reuse the intuition that $p(1-p)$ is the variance of a Bernoulli distribution; in a continuous domain, an analogous metric would be the variance of rewards obtained by playing the same level multiple times. Furthermore, our implementation of SFL is in JAX but the method is general. However, one must take the cost of SFL’s additional environment rollouts into account when considering implementing our algorithm; we chose JAX because its speed and parallelisation significantly alleviates this constraint. Next, while most current SoTA UED methods, including SFL, randomly generate and curate levels, this approach may become infeasible when the environment space is vast, as random generation may have a very low likelihood of generating valid levels. Finally, while JaxNav does have deterministic dynamics, Fan et al. [48] successfully transferred an RL-based multi-robot navigation policy from a simulator of identical fidelity to the real world, suggesting this should be equally possible with JaxNav.
@@ 10 Conclusion
In this paper, we investigate the scoring functions used by current regret-based UED methods and analyse whether they can accurately approximate regret. We find that this is not the case and that these prioritisation metrics instead correlate with success rate, leading to a large amount of experience not contributing to learning an improved policy. Inspired by this analysis, we develop a method based on an intuitive notion of learnability and find that this improves the robustness of the final policies. We also introduce a new robustness-measuring evaluation protocol, reporting a risk measure on performance over the $\alpha\%$ worst-case (but solvable) levels for each method. We hope that our findings inspire future work on more general, domain-agnostic scoring functions, and we open-source all of our code to facilitate this process. Ultimately, we believe this work is a stepping stone towards bridging the gap between popular testbeds for UED and real-world applications.
Acknowledgements
This work received funding from the EPSRC Programme Grant “From Sensing to Collaboration” (EP/V000748/1). JF is partially funded by the UKI grant EP/Y028481/1 (originally selected for funding by the ERC). JF is also supported by the JPMC Research Award and the Amazon Research Award. MB is funded by the Rhodes Trust.
@@ Appendix F Extended analysis of UED Score Functions
An extension of Section 4.2, Figure 12 illustrates the correlation between mean reward and the two most popular regret scores, MaxMC and PVL. These graphs illustrate that the trends seen for success rate also hold for episodic reward, this is expected for our environment as the two are strongly correlated. In Figure 13 we conduct the same analysis for the L1 Loss Score [9], defined as:
[…]